The ratio of the area of two circle is 16/9. What is the ratio of their circumference?
a. 4/3 b. 16/9 c. 32/18 d. 256/81

Answer by edjones(8007)

pi*r^2/pi*r^2=16/9 pi cancels out. area/area
r/r=4/3 sqrt of each side
2pi*r/2p*r=4/3 [2pi cancels out] circumf/circumf
.
Ed